Calculate the kinetic energy required to accelerate a neutron from a rest position to 0.99990 If the mass of the neutron is 1.68*10^ -27 kg


1
Expert's answer
2021-04-19T17:07:29-0400

The relativistic kinetic energy is given as follows:


"K = mc^2\left( \dfrac{1}{\sqrt{1-\dfrac{v^2}{c^2}}} - 1 \right)"

where "m = 1.68\times10^{-27}kg" is the mass of the neutron, "c = 3\times 10^8 m/s" is the speed of light, "v = 0.99990c" is the speed of the neutron. Thus, obtain:


"K = 1.68\times10^{-27}\cdot (3\times 10^8)^2\left( \dfrac{1}{\sqrt{1-\dfrac{(0.99990c)^2}{c^2}}} - 1 \right) \approx 1.05\times 10^{-8}J"

Answer. "1.05\times 10^{-8}J".